OpenStudy (anonymous):

I'm in College Physics and I'm dealing with motion, acceleration, velocity and time intervals. We were given 3 formula and I don't understand them, can someone explain what its for and help me with an example?

OpenStudy (theeric):

I'm not entirely up to par at the moment, but I'll be happy to try to help! So, what's most important to finding out which equation to use is know what variables are in each! So what exactly would you like to know? I'm sure there are quite a few people who can help on OpenStudy!

OpenStudy (anonymous):

v^2-u^2=2as s=ut+.5at^2 v=u+at are these formula given??

OpenStudy (anonymous):

Well they are word problems and I'm able to get the information for the most part. An example is ...Suppose a car is moving with constant velocity along a straight road. Its position was x1 = 16m at t1 = 0.0 s and is x2 = 31m at t2 = 3.0s what is the position at 1.7 second? I did not use the formulas but used logic to solve step 1. the formulas are different.

OpenStudy (anonymous):

(Vx)f=(Vx)i+ax*delta T also Xf=xi+(Vx)i *delta T+1/2ax(delta T)^2 and (vx)f^2=(Vx)i^2+2ax*delta x

OpenStudy (theeric):

Hi! I'm going to make your equations a little easier to read. \[v_{x,f}=v_{x,i}+a_x\Delta t\] \[x_f=x_i+v_{x,i} \Delta t+\frac{1}{2}a_x(\Delta t)^2\] \[v_{x,f}^2=v_{x,i}^2+2a_x\Delta x\] These are all common physics equations! The first one is as simple as the definition of acceleration, rearranged. That is,\[v_{x,f}=v_{x,i}+a_x\Delta\\\qquad\Downarrow\\v_{x,f}-v_{x,i}=a_x\Delta t\\\qquad\Downarrow\\\frac{v_{x,f}-v_{x,i}}{\Delta t}=a_x\]That is the change in velocity per a change in time! The other equations might have required calculus to derive, but the second one is sort of understandable. \[\begin{matrix}x_f&=&x_i&+&v_{x,i} \Delta t&+&\frac{1}{2}a_x(\Delta t)^2\\\uparrow&&\uparrow&&\uparrow&&\uparrow\\1&&2&&3&&4\end{matrix}\]Term 1: final position Term 2: initial position Term 3: change in position from the initial velocity Term 4: change in position from the acceleration Term 1: the initial position plus the total change in position That's an okay way to think of it, just in that you need all of those component for the equation. For all of these, memorization will be handy, and hopefully the first one is more easy to rediscover yourself.

OpenStudy (theeric):

For your example problem, I would use the definition of velocity: \(\Large v =\frac{\Delta x}{\Delta t}=\frac{x_2-x_1}{t_2-t_1}\) And then look at the time \(t_3=1.7\ [s]\) with the equation \(x_f=x_i + v\Delta t\). In case you want to tie this to your equations, look to your second equation and let \(a=0\). It might be weird to be working with \(x_f\) and \(x_i\) when you don't see those variables, but you put your variables into those spots, and \(\Delta t\) is the time difference from \(x_i\) to \(x_f\).

OpenStudy (masumanwar):

Motion is produced when any object move and displaced in space with time and physics relates this two measurable quantity in to some other measurable quantity as velocity is a vector quantity defined as V=ds/dt and acceleration as a=dv/dt where ds,dv,dt,are the small change in those quantity measured in meter,meter per secon, and second